# FRoundation: Are Foundation Models Ready for Face Recognition?

### v1

@@ Abstract
keywords:
Face recognition , Foundation models , Computer vision , Face and gesture analyses , Human analyses , fine-tuning[inst1]organization=Fraunhofer Institute for Computer Graphics Research, city=Darmstadt, postcode=64283, country=Germany
[inst2]organization=Department of Computer Science, TU Darmstadt, city=Darmstadt, postcode=64283, state=Hessen, country=Germany
@@ 1 Introduction
Significant progress has been made in developing foundation models trained on extensive and diverse datasets recently [1, 2, 3]. Once these models are trained, they can be adapted to a wide array of downstream tasks, providing a versatile basis. Inspired by the success of large language models (LLM) [4, 5, 6, 7, 8], similar large-scale foundation models have been explored for various perception tasks [9, 1, 2, 3]. Often based on the Vision Transformer (ViT) architecture [10], which has been shown to match or exceed the performance of traditional methods in large-scale image classification tasks, visual foundation models are becoming increasingly popular due to their strong generalization capabilities when trained on large datasets. Although foundation models hold considerable promise for a wide range of applications, their adaptation for face recognition (FR) has not been explored in previous research, which drives the motivation for this study.
To explore the potential of foundation models in FR, we evaluate the performance of different versions of two widely used foundation models, namely DINOv2 [1] and CLIP [9]. Although foundation models demonstrate strong generalization capabilities, our experiments show that they perform poorly compared to state-of-the-art (SOTA) FR models [11, 12, 13, 14] on various benchmarks. To enhance their effectiveness for the downstream task of FR, we propose to fine-tune the considered foundation models using low-rank adaptation (LoRA) [15]. LoRA integrates trainable low-rank decomposition matrices into each transformer block while keeping the pre-trained model weights frozen. After fine-tuning on dedicated datasets, the LoRA layers adapt to the downstream task of FR, resulting in a significant increase in performance. For example, the smallest released pretrained models of DINOv2 and CLIP achieve average verification accuracies of 64.70% and 82.64%, respectively. After fine-tuning on Casia-Webface, their accuracies increase to 90.94% and 92.13%, respectively.
In addition to examining the performance of fine-tuned foundation models, we compare them to a ViT trained from scratch on different subset sizes of data. This comparison validates the advantages of using foundation models over training an entire model based on data availability. Finally, we investigate the bias in FR and demonstrate that it tends to be slightly higher when fine-tuning foundation models compared to training from scratch.
@@ 2 Related Work
Foundation models are defined by a substantial number of trainable parameters and are pre-trained on a large and diverse dataset, enabling them to adapt to a wide range of downstream tasks across various domains [16]. Vision foundation models are commonly structured around the use of ViTs [10] and rely on self-supervised learning (SSL) [17]. SSL is a technique that trains models to learn representations from unlabeled data and is essential for training ViTs, which tend to perform poorly on small datasets [18]. Several high-performing vision foundation models specialized in various tasks have been released in recent years. Building on the success of the Segment Anything Model (SAM) [19], which introduced a foundation model for object segmentation, SAM2 [20] presents a model for real-time, promptable object segmentation in images and videos, achieving state-of-the-art performance. CLIP [9] learns visual concepts from natural language supervision and can be applied to any visual classification by providing the names of the visual categories to be recognized. DINOv2 [1] foundation models generate universal features suitable for both image-level visual tasks, such as image classification and video understanding, as well as pixel-level visual tasks, including depth estimation and semantic segmentation.
@@
Vision foundation models are characterized by their generalization ability due to massive training data, but they tend to show poor performance when applied to domain-specific settings [21]. To adapt them to a downstream task, multiple approaches were considered in the literature for this adaption. An example of that is the AdaptFormer [22] that replaces the MLP block in the transformer encoder with two identical MLP branches, where one mirrors the original network and the other introduces a lightweight module for task-specific fine-tuning, demonstrating significant improvements compared to fully fine-tuned models on downstream tasks. Another example is the ViT-Adapter [23] that proposes a technique to adapt plain ViTs for dense prediction tasks by injecting image-related inductive biases into the ViT and reconstructing fine-grained multi-scale features, yielding SOTA results on COCO test-dev. Another approach is to insert trainable rank decomposition matrices, called LoRA [15] layers, while freezing the pre-trained model weights. In this work, we choose LoRA as our foundation model adapter, as recent studies [24, 25, 26, 27, 21] highlight the significant potential of LoRA for this purpose. For instance, integrating LoRA layers into DINOv2 has been successfully applied in the medical domain for two distinct tasks: capsule endoscopy diagnosis [24] and depth estimation in endoscopic surgery [25], both of which have demonstrated superior performance in their respective fields. In another work on adapting foundation models for multiple plant phenotyping [26], Chen et al. demonstrated that LoRA consistently outperforms decoder tuning in leaf counting and disease classification, with their method achieving high performance in both tasks, approaching the results of SOTA models. Zanella et al. [27] explored few-shot adaptation of Vision Language Models (VLMs) using LoRA, showing that their CLIP-based method not only achieves substantial improvements but also reduces training times.
@@
The application of foundation models in biometrics is still very limited, with only few recent works started investigating their potential. For example, Iris-SAM [28], which is based on the foundation model SAM [19], fine-tunes it on ocular images for iris segmentation, achieving an average segmentation accuracy that surpasses the best baseline by a substantial margin of 10% on the ND-IRIS-0405 dataset. Arc2Face [29] is an identity-conditioned face foundation model that generates photo-realistic images based on the ArcFace embedding of a person. To showcase the performance of the generated data, they train a FR model on synthetic images from their model, achieving superior performance compared to existing synthetic datasets [30]. Recognizing the immense potential of foundation models across diverse tasks, this study breaks new ground by exploring their adaptation for FR, a path, to the best of our knowledge, unexplored until now.
@@ 3 Methodology
This section presents our approach FRoundation, for optimizing vision foundation models for FR. This section starts by providing details on the selected baseline foundation models, CLIP [9] and DINOv2 [1]. Then, we provide details on the adaptation mechanism used to optimize foundation models for downstream tasks. Finally, we conclude by describing the extension of pretrained foundation models with LoRA for FR.
@@ 3.1.1 DINOv2
DINOv2 [1] is a self-supervised image encoder trained on a large curated dataset, namely the LVD-142M dataset. The dataset was created as part of this initiative, using an automated pipeline to assemble a dedicated, diverse, and curated collection of images.The model network architecture follows a student-teacher framework based on ViT [10] that learns features at the image and patch levels by combining DINO [31] and iBOT [32]losses. The image-level objective, inspired by DINO, is derived from the cross-entropy loss between features extracted from the student and teacher networks. These features are taken from the ViT class token and are based on different crops of the same image. For the patch-level objective, random input patches are masked and sent to the student, while remaining visible to the teacher. The student’s iBOT head processes the masked tokens while the teacher’s iBOT head processes the corresponding visible tokens, leading to the calculation of the loss term. Additionally, several contributions were made to accelerate and stabilize large-scale training. As a result, a ViT model with 1B parameters was trained and distilled into a series of smaller models that surpass the best available general-purpose features on most of the benchmarks at image and pixel levels [1], making it a top candidate as a foundation model for FR.
@@ 3.3 FRoundation
Utilizing pretrained foundation model for FR leads to suboptimal results, as will be presented in Section 5.1. Thus, we propose to fine-tune the considered foundation models using LoRA. This requires extending the pretrained ViT models with LoRA layers before fine-tuning them on dedicated datasets. It is possible to apply LoRA to the , , , and projection layers, which refer respectively to the query, key, value, and output projection matrices in the self-attention module of ViT architecture [10]. Following [15], we adapt the weight matrices of q and v only, as it was shown that adapting their respective weights and yields the best results on different downstream tasks [15].
@@
The output of the projection layer is then passed to the MLP, completing the execution of a single transformer block. transformer layers are used to transform the image tokens into feature representations where denotes the output of the -th transformer block. Each -th transformer layer processes the output vectors of the previous layer. The final hidden state of the CLS token is employed as feature representation. To optimize the considered foundation models for FR, we proposed to fine-tune the models with the widely adapted margin-penalty softmax loss for FR [12, 11]. Specifically, we extended the architecture with an additional multi-class classification layer and utilized CosFace loss [12] as a margin-penalty softmax loss.
@@ 4 Experimental setups
To explore the capabilities of foundation models for FR, we assess their performance using the following benchmarks: Labeled Faces in the Wild (LFW) [36], Celebrities in Frontal-Profile in the Wild (CFP-FP) [37], AgeDB30 [38], Cross-age LFW (CA-LFW) [39], CrossPose LFW (CP-LFW) [40]. The results are reported on each benchmark as verification accuracies in (%) and following their official evaluation protocol. In addition, we evaluated on large-scale evaluation benchmarks, IARPA Janus Benchmark-B (IJB-B) [41], and IARPA Janus Benchmark–C (IJB-C) [42]. For IJB-C and IJB-B, we used the official 1:1 mixed verification protocol and reported the verification performance as true acceptance rates (TAR) at false acceptance rates (FAR) of 1e-3, 1e-4, and 1e-5. All these benchmarks were chosen as they are widely used to benchmark the latest FR developments and provide a comprehensive variation of use-cases [11, 12, 13, 14].
@@
We utilize the CosFace [12] loss function to train all the presented models in this paper with a margin penalty of and scale factor of , following [12], as well as other works analyzing different building blocks of FR [48]. During the fine-tuning, we used AdamW [49] optimizer with a weight decay of 0.05. We train for 40 epochs (on CaisaWebFace) and for 30 epochs (on MS1MV2 and WebFace4M), with a batch size set to 512 [11]. The initial learning rate is set to 0.0001 and is updated using a Cosine Learning Rate scheduler [50]. Additionally, the LoRA rank is set to 16 for all applicable experiments. The images are resized to pixels to adapt to the image resolution initially used by DINOv2 and CLIP during training. For data augmentation, we apply horizontal flipping and RandAug [51] with a number of operations and a magnitude of , following [52]. We additionally trained 12 instances of ViT, ViT-S and ViT-L, from scratch on different subsets of CASIA-WebFace as well as on MS1MV2 and WebFace4M. These models are noted as Baseline.
@@ 5.3 Large-scale Fine-tuning
To conclude, for the case where only a relatively small training dataset is accessible, finetuning pretrained foundation models can achieve higher recognition accuracy than training the same model architecture from scratch. In the case that one has access to large training datasets, training a model from scratch for FR can achieve very competitive results to finetuning foundation models. However, this, next to the technical and legal limitations of collecting or maintaining the data, comes with a high training time cost. For example, training ViT-L from scratch on MS1MV2 using the settings described in Section 4 requires around 70 GPU hours on MS1MV2, in comparison to around 40 GPU hours for finetuning DINOv2 with LoRA on 8 Nvidia A100 SXM4 40 GB GPUs.
@@ 6 Conclusion
This paper was the first to propose and investigate the use of foundation models for the task of FR. We additionally propose the adaption of the foundation models to this specific task under various levels of data availability. Our experiments on multiple foundation models, training datasets, and a wide range set of evaluation benchmarks led to interesting conclusions summarized in the following. The studied pretrained foundation models, used as feature extractors without fine-tuning, demonstrated relatively acceptable (far from random) accuracy on less challenging face verification benchmarks like LFW. However, they under-performed on more challenging benchmarks such as AgeDB30, CALFW, IJB-B, and IJB-C. Our adaption of foundation models to FR showed that fine-tuning these models on even small subsets of CASIA-WebFace, e.g. 1k identities, significantly boosts their verification accuracy across benchmarks, outperforming models trained from scratch on similar data subsets. Additionally, increasing the dataset size (dataset width) or the number of images per identity (dataset depth) enhances performance, demonstrating the scalability of foundation models with diverse data availability. The results additionally showed that, with a limited training dataset like CASIA-WebFace, fine-tuning pretrained foundation models outperforms training models from scratch in recognition accuracy. However, when large datasets (MS1MV2 or WebFace4M) are available, training from scratch yields competitive performance, though it incurs a significantly higher training computational cost compared to fine-tuning. Our bias evaluation results indicate that fine-tuning foundation models enhance recognition performance across demographics but introduces slightly more bias compared to baseline models trained from scratch, especially on smaller fine-tuning datasets. Models trained from scratch on larger datasets (MS1MV2 and WebFace4M) achieved superior recognition accuracy and exhibited lower bias. The outcomes of this work set the stage for broader adoption of foundation models as a basis for biometric recognition, particularly in scenarios with limited data availability, being aware of the technical and legal constraints on biometric data collection and management.
Acknowledgments
This research work has been funded by the German Federal Ministry of Education and Research and the Hessen State Ministry for Higher Education, Research and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE.

### v2

@@ 1 Introduction
Significant progress has been made in developing foundation models trained on extensive and diverse datasets recently [33, 2, 23]. Once these models are trained, they can be adapted to a wide array of downstream tasks, providing a versatile basis. Inspired by the success of large language models (LLM) [18, 8, 39, 13, 38], similar large-scale foundation models have been explored for various perception tasks [35, 33, 2, 23]. Often based on the ViT architecture [19], which has been shown to match or exceed the performance of traditional methods in large-scale image classification tasks, visual foundation models are becoming increasingly popular due to their strong generalization capabilities when trained on large datasets. Although foundation models hold considerable promise for a wide range of applications, their adaptation for face recognition has not been explored in previous research, which drives the motivation for this study.
To explore the potential of foundation models in face recognition, we evaluate the performance of different versions of two widely used foundation models, namely DINOv2 [33] and CLIP [35]. Although foundation models demonstrate strong generalization capabilities, our experiments show that they perform poorly compared to state-of-the-art (SOTA) face recognition models [17, 42, 4, 16] on various benchmarks. To enhance their effectiveness for the downstream task of face recognition, we propose to fine-tune the considered foundation models using low-rank adaptation (LoRA) [24]. LoRA integrates trainable low-rank decomposition matrices into each transformer block while keeping the pre-trained model weights frozen. After fine-tuning on dedicated datasets, the LoRA layers adapt to the downstream task of face recognition, resulting in a significant increase in performance. For example, the smallest released pretrained models of DINOv2 and CLIP achieve average verification accuracies of 64.70% and 82.64%, respectively. After fine-tuning on Casia-Webface, their accuracies increase to 90.94% and 92.13%, respectively.
In addition to examining the performance of fine-tuned foundation models, we compare them to a Vision Transformer (ViT) trained from scratch on different subset sizes of data. This comparison validates the advantages of using foundation models over training an entire model based on data availability. Finally, we investigate the bias in face recognition and demonstrate that it tends to be slightly higher when fine-tuning foundation models compared to training from scratch.
@@ 2 Related Work
Foundation models are defined by a substantial number of trainable parameters and are pre-trained on a large and diverse dataset, enabling them to adapt to a wide range of downstream tasks across various domains [3]. Vision foundation models are commonly structured around the use of Vision Transformers (ViTs) [19] and rely on self-supervised learning (SSL) [21]. SSL is a technique that trains models to learn representations from unlabeled data and is essential for training ViTs, which tend to perform poorly on small datasets [54]. Several high-performing vision foundation models specialized in various tasks have been released in recent years. Building on the success of the Segment Anything Model (SAM) [28], which introduced a foundation model for object segmentation, SAM2 [36] presents a model for real-time, promptable object segmentation in images and videos, achieving state-of-the-art performance. CLIP [35] learns visual concepts from natural language supervision and can be applied to any visual classification by providing the names of the visual categories to be recognized. DINOv2 [33] foundation models generate universal features suitable for both image-level visual tasks, such as image classification and video understanding, as well as pixel-level visual tasks, including depth estimation and semantic segmentation.
@@
Vision foundation models are characterized by their generalization ability due to massive training data, but they tend to show poor performance when applied to domain-specific settings [41]. To adapt them to a downstream task, multiple approaches were considered in the literature for this adaption. An example of that is the AdaptFormer [11] that replaces the MLP block in the transformer encoder with two identical MLP branches, where one mirrors the original network and the other introduces a lightweight module for task-specific fine-tuning, demonstrating significant improvements compared to fully fine-tuned models on downstream tasks. Another example is the ViT-Adapter [12] that proposes a technique to adapt plain ViTs for dense prediction tasks by injecting image-related inductive biases into the ViT and reconstructing fine-grained multi-scale features, yielding SOTA results on COCO test-dev. Another approach is to insert trainable rank decomposition matrices, called Low-rank adaptation (LoRA) [24] layers, while freezing the pre-trained model weights. In this work, we choose LoRA as our foundation model adapter, as recent studies [49, 15, 10, 48, 41] highlight the significant potential of LoRA for this purpose. For instance, integrating LoRA layers into DINOv2 has been successfully applied in the medical domain for two distinct tasks: capsule endoscopy diagnosis [49] and depth estimation in endoscopic surgery [15], both of which have demonstrated superior performance in their respective fields. In another work on adapting foundation models for multiple plant phenotyping [10], Chen et al. demonstrated that LoRA consistently outperforms decoder tuning in leaf counting and disease classification, with their method achieving high performance in both tasks, approaching the results of SOTA models. Zanella et al. [48] explored few-shot adaptation of Vision Language Models (VLMs) using LoRA, showing that their CLIP-based method not only achieves substantial improvements but also reduces training times.
@@
The application of foundation models in biometrics is still very limited, with only few recent works started investigating their potential. For example, Iris-SAM [20], which is based on the foundation model SAM [28], fine-tunes it on ocular images for iris segmentation, achieving an average segmentation accuracy that surpasses the best baseline by a substantial margin of 10% on the ND-IRIS-0405 dataset. Arc2Face [34] is an identity-conditioned face foundation model that generates photo-realistic images based on the ArcFace embedding of a person. To showcase the performance of the generated data, they train a FR model on synthetic images from their model, achieving superior performance compared to existing synthetic datasets [7]. Recognizing the immense potential of foundation models across diverse tasks, this study breaks new ground by exploring their adaptation for face recognition, a path, to the best of our knowledge, unexplored until now.
@@ 3 Methodology
This section presents our approach FRoundation, for optimizing vision foundation models for face recognition. This section starts by providing details on the selected baseline foundation models, CLIP [35] and DINOv2 [33]. Then, we provide details on the adaptation mechanism used to optimize foundation models for downstream tasks. Finally, we conclude by describing the extension of pretrained foundation models with LoRA for face recognition.
@@ 3.1.1 DINOv2
DINOv2 [33] is a self-supervised image encoder trained on a large curated dataset, namely the LVD-142M dataset. The dataset was created as part of this initiative, using an automated pipeline to assemble a dedicated, diverse, and curated collection of images.The model network architecture follows a student-teacher framework based on vision transformers (ViT) [19] that learns features at the image and patch levels by combining DINO [9] and iBOT [53]losses. The image-level objective, inspired by DINO, is derived from the cross-entropy loss between features extracted from the student and teacher networks. These features are taken from the ViT class token and are based on different crops of the same image. For the patch-level objective, random input patches are masked and sent to the student, while remaining visible to the teacher. The student’s iBOT head processes the masked tokens while the teacher’s iBOT head processes the corresponding visible tokens, leading to the calculation of the loss term. Additionally, several contributions were made to accelerate and stabilize large-scale training. As a result, a ViT model with 1B parameters was trained and distilled into a series of smaller models that surpass the best available general-purpose features on most of the benchmarks at image and pixel levels [33], making it a top candidate as a foundation model for FR.
@@ 3.3 FRoundation
Utilizing pretrained foundation model for face recognition leads to suboptimal results, as will be presented in Section 5.1. Thus, we propose to fine-tune the considered foundation models using LoRA. This requires extending the pretrained ViT models with LoRA layers before fine-tuning them on dedicated datasets. It is possible to apply LoRA to the , , , and projection layers, which refer respectively to the query, key, value, and output projection matrices in the self-attention module of ViT architecture [19]. Following [24], we adapt the weight matrices of q and v only, as it was shown that adapting their respective weights and yields the best results on different downstream tasks [24].
@@
The output of the projection layer is then passed to the MLP, completing the execution of a single transformer block. transformer layers are used to transform the image tokens into feature representations where denotes the output of the -th transformer block. Each -th transformer layer processes the output vectors of the previous layer. The final hidden state of the CLS token is employed as feature representation. To optimize the considered foundation models for FR, we proposed to fine-tune the models with the widely adapted margin-penalty softmax loss for face recognition [42, 17]. Specifically, we extended the architecture with an additional multi-class classification layer and utilized CosFace loss [42] as a margin-penalty softmax loss.
@@ 4 Experimental setups
To explore the capabilities of foundation models for face recognition, we assess their performance using the following benchmarks: Labeled Faces in the Wild (LFW) [25], Celebrities in Frontal-Profile in the Wild (CFP-FP) [37], AgeDB30 [32], Cross-age LFW (CA-LFW) [52], CrossPose LFW (CP-LFW) [51]. The results are reported on each benchmark as verification accuracies in (%) and following their official evaluation protocol. In addition, we evaluated on large-scale evaluation benchmarks, IARPA Janus Benchmark-B (IJB-B) [46], and IARPA Janus Benchmark–C (IJB-C) [31]. For IJB-C and IJB-B, we used the official 1:1 mixed verification protocol and reported the verification performance as true acceptance rates (TAR) at false acceptance rates (FAR) of 1e-3, 1e-4, and 1e-5. All these benchmarks were chosen as they are widely used to benchmark the latest FR developments and provide a comprehensive variation of use-cases [17, 42, 4, 16].
@@
We utilize the CosFace [42] loss function to train all the presented models in this paper with a margin penalty of and scale factor of , following [42], as well as other works analyzing different building blocks of face recognition [5]. During the fine-tuning, we used AdamW [30] optimizer with a weight decay of 0.05. We train for 40 epochs (on CaisaWebFace) and for 30 epochs (on MS1MV2 and WebFace4M), with a batch size set to 512 [17]. The initial learning rate is set to 0.0001 and is updated using a Cosine Learning Rate scheduler [29]. Additionally, the LoRA rank is set to 16 for all applicable experiments. The images are resized to pixels to adapt to the image resolution initially used by DINOv2 and CLIP during training. For data augmentation, we apply horizontal flipping and RandAug [14] with a number of operations and a magnitude of , following [6]. We additionally trained 12 instances of Vit, Vit-S and Vit-L, from scratch on different subsets of CASIA-WebFace as well as on MS1MV2 and WebFace4M. These models are noted as Baseline.
@@ 5.3 Large-scale Fine-tuning
To conclude, for the case where only a relatively small training dataset is accessible, finetuning pretrained foundation models can achieve higher recognition accuracy than training the same model architecture from scratch. In the case that one has access to large training datasets, training a model from scratch for face recognition can achieve very competitive results to finetuning foundation models. However, this, next to the technical and legal limitations of collecting or maintaining the data, comes with a high training time cost. For example, training Vit-L from scratch on MS1MV2 using the settings described in Section 4 requires around 70 GPU hours on MS1MV2, in comparison to around 40 GPU hours for finetuning DINOv2 with LoRA on 8 Nvidia A100 SXM4 40 GB GPUs.
@@ 6 Conclusion
This paper was the first to propose and investigate the use of foundation models for the task of face recognition. We additionally propose the adaption of the foundation models to this specific task under various levels of data availability. Our experiments on multiple foundation models, training datasets, and a wide range set of evaluation benchmarks led to interesting conclusions summarized in the following. The studied pretrained foundation models, used as feature extractors without fine-tuning, demonstrated relatively acceptable (far from random) accuracy on less challenging face verification benchmarks like LFW. However, they under-performed on more challenging benchmarks such as AgeDB30, CALFW, IJB-B, and IJB-C. Our adaption of foundation models to face recognition showed that fine-tuning these models on even small subsets of CASIA-WebFace, e.g. 1k identities, significantly boosts their verification accuracy across benchmarks, outperforming models trained from scratch on similar data subsets. Additionally, increasing the dataset size (dataset width) or the number of images per identity (dataset depth) enhances performance, demonstrating the scalability of foundation models with diverse data availability. The results additionally showed that, with a limited training dataset like CASIA-WebFace, fine-tuning pretrained foundation models outperforms training models from scratch in recognition accuracy. However, when large datasets (MS1MV2 or WebFace4M) are available, training from scratch yields competitive performance, though it incurs a significantly higher training computational cost compared to fine-tuning. Our bias evaluation results indicate that fine-tuning foundation models enhance recognition performance across demographics but introduces slightly more bias compared to baseline models trained from scratch, especially on smaller fine-tuning datasets. Models trained from scratch on larger datasets (MS1MV2 and WebFace4M) achieved superior recognition accuracy and exhibited lower bias. The outcomes of this work set the stage for broader adoption of foundation models as a basis for biometric recognition, particularly in scenarios with limited data availability, being aware of the technical and legal constraints on biometric data collection and management.